A regression was run to determine if there is a relationship between hours of TV watched per day (x) and number of situps a person can do (y).

The results of the regression were:

y=ax+b
a=-0.949
b=30.119
r2=0.597529
r=-0.773



Assume the correlation is significant, and use this to predict the number of situps a person who watches 10 hours of TV can do (to one decimal place)
Answer #1
 We have the regression equation y = ax+b

= - 0.949 (x) + 30.119

When x=10

 y = -0.949(10) + 30.119

= 20.629

The predicted number of situps a person who watches 10 hours of TV can do is = 20.6